### v1

@@ 1 Introduction
A key axiom for the Nash solution is arrow1959rational’s choice axiom, a set-valued version of Nash’s independence of irrelevant alternatives (cf. kaneko1980extension; xu2006alternative). This axiom states that if the feasible set expands and the rule still chooses some alternatives from the original set, then all originally chosen alternatives must still be selected (i.e., for two problems $S,S^{\prime}$ such that $S\subseteq S^{\prime}$, $F(S^{\prime})\cap S\neq\emptyset$ implies $F(S)=F(S^{\prime})\cap S$). This property is closely related to the rationalizability of multi-valued bargaining solutions or choice correspondences via welfare functions or orderings (ok1999revealed).
@@
This observation raises a natural question: What rationality axioms should we impose to construct collective choice rules, and what class of rules satisfies them? To address this issue, we propose two axioms by weakening Arrow’s choice axiom. The first axiom modifies the dual Chernoff axiom while retaining the Chernoff axiom. Recall that the original dual Chernoff axiom implies that if an outcome is chosen in both $S$ and $T$, it should also be chosen in $S\cup T$. Our axiom also requires the converse: If an outcome is chosen in $S\cup T$ and is feasible in both $S$ and $T$, it must be chosen in both. We call this axiom the weak Arrow axiom because it coincides with the combination of the Chernoff axiom and the weak dual Chernoff axiom known in the literature. The second axiom postulates that all outcomes chosen in the original problem should be selected in an expanded problem if no outcome that becomes feasible owing to the expansion is chosen. This prevents the disappearance of all chosen outcomes upon irrelevant expansions; we therefore call this axiom independence of irrelevant expansions. We investigate implications of these axioms under standard axioms such as Pareto optimality, scale invariance, and continuity.
@@ 2 The bargaining problem
We introduce several notations for convenience. A permutation is a bijection on $N$. Let $\Pi$ be the set of permutations. For $x\in\mathbb{R}^{n}$ and $\pi\in\Pi$, define $x_{\pi}$ as the vector $(x_{\pi(1)}.x_{\pi(2)},\ldots,x_{\pi(n)})\in\mathbb{R}^{n}$. A set $S\subseteq\mathbb{R}^{n}$ is said to be symmetric if $S=\{x_{\pi}\mid x\in S\}$ for all $\pi\in\Pi$. For $a,x\in\mathbb{R}^{n}$, let $a\ast x=(a_{1}x_{1},a_{2}x_{2},\ldots,a_{n}x_{n})$. Similarly, for $a\in\mathbb{R}^{n}$ and $S\subseteq\mathbb{R}^{n}$, let $a\ast S=\{a\ast x\mid x\in S\}$.
@@
A special case with equal weight (i.e., $w_{i}=w_{j}$ for all $i,j\in N$) is the Nash solution $F^{\text{N}}$. That is, for all $S\in\mathcal{B}$,
@@ 3 Rationalizabiliy
Based on this decomposition, the requirements of Arrow axiom can be understood as follows. Suppose that a choice set $S$ expands to $S^{\prime}$ and the rule still chooses some outcomes from the original problem $S$ (that is, $F(S^{\prime})\cap S\neq\emptyset$). Then, the first condition requires that the choice from the expanded problem $S^{\prime}$ that was available in the original problem should only be taken from the original chosen outcomes (i.e., Chernoff Axiom); and the second condition requires that all the chosen outcomes in the original problem $S$ must remain chosen (i.e., dual Chernoff Axiom).
@@
For all $S,T\in\mathcal{B}$, $F(S)\cap F(T)\subseteq F(S\cup T)$. [^4] This axiom is sometimes written as follows: For any $\{S^{k}\}_{k=1}^{K}\subseteq\mathcal{B}$, $\bigcap_{k=1}F(S^{k})\subseteq F(\bigcup_{k=1}^{K}S^{k})$. By induction, the equivalence of both axioms can be shown, so that we can use the present axiom without loss of generality.
@@
For all $S,S^{\prime}\in\mathcal{B}$ with $S\subseteq S^{\prime}$, if $F(S^{\prime})\subseteq S$, then $F(S)\subseteq F(S^{\prime})\cap S$.
Weak dual Chernoff axiom is already known in the literature on social choice theory. It requires that, for any problem $S$ and $T$, the commonly chosen outcomes in both problems should also be chosen in the combined problem $S\cup T$. This axiom is indeed weaker than dual Chernoff axiom because the latter implies both $F(S)\subseteq F(S\cup T)$ and $F(T)\subseteq F(S\cup T)$. Note that this axiom can accommodate the intuitive decision in the above scenarios. IIE is a new axiom about problem expansions. It requires that if the expansion is irrelevant in the sense that all the outcomes chosen in the expanded problem $S^{\prime}$ can be taken in the original problem $S$, then the outcomes chosen in $S$ should be chosen in $S^{\prime}$ again. This axiom avoids the above criticism of Arrow axiom by strengthening the prerequisite of dual Chernoff axiom.
As weaker rationality axioms for bargaining solutions than Arrow axiom, we adopt Chernoff axiom, weak dual Chernoff axiom, and IIE, and examine their implications under the aforementioned standard axioms. We first show that the implications of the combination of Chernoff axiom and weak dual Chernoff axiom are summarized as follows:
@@
Remark. In the finite-alternative case, sen1971choice shows the counterpart result of (2) in Proposition 2, using Chernoff axiom, weak dual Chernoff axiom, and so-called Sen’s $\delta$. [^5] In our model, it can be said that $F$ satisfies Sen’s $\delta$ if, for all $S,S^{\prime}\in\mathcal{B}$ with $S\subseteq S^{\prime}$, if $\{x,y\}\subseteq F(S)$, then $\{x\}\neq F(S^{\prime})$. A stronger version of the axiom can be stated that, for all $S,S^{\prime}\in\mathcal{B}$ with $S\subseteq S^{\prime}$, if $A\subseteq F(S)$, then $B\neq F(S^{\prime})$ for all $B\subsetneq A$. Even under this stronger version, our result cannot be established. Hence, technically speaking, this result is a generalization of Sen’s result to an infinite-alternative case by using a different axiom–IIE–with efficiency and continuity to address the additional complexity caused by infinitely many alternatives.
@@ 4 The coarse Nash solutions
Putting differently, this proposition states that for each coarse Nash solution $F^{A}$, there exists $w\in\Delta_{N}$ such that for all $x,y\in\mathbb{R}^{n}_{++}$,
@@ 5 Related literature
It should be noted that we employ a domain different from nash1950bargaining’s (nash1950bargaining) original formalization. In the most standard framework of the bargaining problem, feasible sets are assumed to be convex, and bargaining solutions are formalized as choice functions. By contrast, we do not assume the convexity of feasible sets and allow solutions to select multiple outcomes. Non-convex bargaining problems have been considered in several studies, such as kaneko1980extension, herrero1989nash, conley1991bargaining, zhou1997nash, ok1999revealed, mariotti2000maximal, xu2006alternative; xu2013rationality; xu2019equitable, and peters2012wpo. For surveys, see also xu2020nonconvex and Thomson2022RED. The convexity has been justified through the assumptions that randomization among agreements is allowed and that each player follows the expected utility theory. In the non-convex problem, we can analyze the situations in which these assumptions do not hold. Note that in such cases, scale invariance still has compelling interpretations, even though this axiom has traditionally been justified within the expected utility framework. For a detailed discussion, see xu2006alternative.
A coarse Nash solution can be regarded as the combination of maximizers of multiple choice criteria, as illustrated in the examples we discussed earlier. The idea of conjoining multiple choice criteria is closely related to tadenuma2002efficiency; tadenuma2005egalitarian and houy2009lexicographic, where the choice correspondence is defined through the sequential application of efficiency and fairness criteria in the context of general resource allocation problems. See also manzini2012choice for a general model of the sequential composition of different choice criteria. In contrast to this sequential approach, our choice correspondence takes into account the intersection of multiple criteria. Such intersections typically involve trade-offs, which often result in incompleteness and render the resulting predictions coarse (e.g., kamaga2018utilitarianism).
@@ A C Proof of Theorem 1
We first show the only-if part. Let $F$ be a solution that satisfies weak Arrow axiom, IIE, efficiency, scale invariance, and continuity. By (2) of Proposition 2, we have known that
@@
Next, we prove the if part. It is easy to prove that the coarse Nash solutions satisfy efficiency. To verify the remained axioms, let $F^{A}$ be some coarse Nash solution.

### v2

@@ 1 Introduction
A key axiom for the Nash solution is arrow1959rational’s (arrow1959rational) choice axiom, a set-valued version of Nash’s independence of irrelevant alternatives (cf. kaneko1980extension; xu2006alternative). This axiom states that if the feasible set expands and the rule still chooses some alternatives from the original set, then all originally chosen alternatives must still be selected (i.e., for two problems $S,S^{\prime}$ such that $S\subseteq S^{\prime}$, $F(S^{\prime})\cap S\neq\emptyset$ implies $F(S)=F(S^{\prime})\cap S$). This property is closely related to the rationalizability of multi-valued bargaining solutions or choice correspondences via welfare functions or orderings (ok1999revealed).
@@
This observation raises a natural question: What rationality axioms should we impose to construct collective choice rules, and what class of rules satisfies them? To address this issue, we introduce two axioms by weakening Arrow’s choice axiom. The first axiom modifies the dual Chernoff axiom while retaining the Chernoff axiom. Recall that the original dual Chernoff axiom implies that if an outcome is chosen in both $S$ and $T$, it should also be chosen in $S\cup T$. Our axiom also requires the converse: If an outcome is chosen in $S\cup T$ and is feasible in both $S$ and $T$, it must be chosen in both. We call this axiom the weak Arrow axiom because it coincides with the combination of the Chernoff axiom and the weak dual Chernoff axiom known in the literature. The second axiom postulates that all outcomes chosen in the original problem should be selected in an expanded problem if no outcome that becomes feasible owing to the expansion is chosen. This prevents the disappearance of all chosen outcomes upon irrelevant expansions; we therefore call this axiom independence of irrelevant expansions. We investigate implications of these axioms under standard axioms such as efficiency, scale invariance, and continuity.
@@ 2 The bargaining problem
We introduce several notations for convenience. A permutation is a bijection on $N$. Let $\Pi$ be the set of permutations. For $x\in\mathbb{R}^{n}$ and $\pi\in\Pi$, define $x_{\pi}$ as the vector $(x_{\pi(1)},x_{\pi(2)},\ldots,x_{\pi(n)})\in\mathbb{R}^{n}$. A set $S\subseteq\mathbb{R}^{n}$ is said to be symmetric if $S=\{x_{\pi}\mid x\in S\}$ for all $\pi\in\Pi$. For $a,x\in\mathbb{R}^{n}$, let $a\ast x=(a_{1}x_{1},a_{2}x_{2},\ldots,a_{n}x_{n})$. Similarly, for $a\in\mathbb{R}^{n}$ and $S\subseteq\mathbb{R}^{n}$, let $a\ast S=\{a\ast x\mid x\in S\}$.
@@
A special case with an equal-weight vector (i.e., $w_{i}=w_{j}$ for all $i,j\in N$) is the Nash solution $F^{\text{N}}$. That is, for all $S\in\mathcal{B}$,
@@ 3 Rationalizabiliy
Based on this decomposition, the requirements of Arrow axiom can be understood as follows. Suppose that a choice set $S$ expands to $S^{\prime}$ and the rule still chooses some outcomes from the original problem $S$ (that is, $F(S^{\prime})\cap S\neq\emptyset$). Then, the first condition requires that the choice from the expanded problem $S^{\prime}$ that was available in the original problem should only be taken from the original chosen outcomes (i.e., Chernoff axiom); and the second condition requires that all the chosen outcomes in the original problem $S$ must remain chosen (i.e., dual Chernoff axiom). [^4] The Chernoff axiom and Dual Chernoff axiom are also called Sen’s $\alpha$ and $\beta$, respectively (sen1970collective).
@@
For all $S,T\in\mathcal{B}$, $F(S)\cap F(T)\subseteq F(S\cup T)$. [^5] This axiom is also called Sen’s $\gamma$ (sen1971choice), which is sometimes written as follows: For any $\{S^{k}\}_{k=1}^{K}\subseteq\mathcal{B}$, $\bigcap_{k=1}F(S^{k})\subseteq F(\bigcup_{k=1}^{K}S^{k})$. By induction, the equivalence of both axioms can be shown, so that we can use the present axiom without loss of generality.
@@
For all $S,S^{\prime}\in\mathcal{B}$ with $S\subseteq S^{\prime}$, if $F(S^{\prime})\subseteq S$, then $F(S)\subseteq F(S^{\prime})$.
Weak dual Chernoff axiom is already known in the literature on social choice theory. It requires that, for any problem $S$ and $T$, the commonly chosen outcomes in both problems should also be chosen in the combined problem $S\cup T$. This axiom is indeed weaker than dual Chernoff axiom because the latter implies both $F(S)\subseteq F(S\cup T)$ and $F(T)\subseteq F(S\cup T)$. Note that this axiom can accommodate the intuitive decision in the above scenarios. IIE is an axiom about problem expansions.666panda1983non calls this axiom Sen’s $\delta^{*}$, which is a stronger version of Sen’s $\delta$ (sen1971choice). cato2018choice calls this axiom weak Nash axiom $\beta$. It requires that if the expansion is irrelevant in the sense that all the outcomes chosen in the expanded problem $S^{\prime}$ can be taken in the original problem $S$, then the outcomes chosen in $S$ should be chosen in $S^{\prime}$ again. This axiom avoids the above criticism of Arrow axiom by strengthening the prerequisite of dual Chernoff axiom.
We adopt Chernoff axiom, weak dual Chernoff axiom, and IIE, and examine their implications under the aforementioned standard axioms. We first show that the implications of the combination of Chernoff axiom and weak dual Chernoff axiom are summarized as follows:
@@
In the finite-alternative case, arrow1959rational established the well-known characterization of rationalizability by Arrow axiom, and tyson2008cognitive provided the corresponding characterization of weak rationalizability by axioms corresponding to weak Arrow axiom and IIE (see also Sen (1971) for another characterization). The first statement in Lemma 1 is a minor modification of ok1999revealed’s (ok1999revealed) extension of Arrow’s result from the finite setting to the non-convex bargaining problem. In parallel, the second statement extends tyson2008cognitive’s (tyson2008cognitive) characterization of weak rationalizability to the non-convex bargaining domain. As in ok1999revealed, the additional axioms of efficiency and continuity are imposed to address the complexity stemming from the difference in domains.
@@ 4 The coarse Nash solutions
To put it differently, this proposition states that for each coarse Nash solution $F^{A}$, there exists $w\in\Delta_{N}$ such that for all $x,y\in\mathbb{R}^{n}_{++}$,
@@ 5 Related literature
It should be noted that we employ a domain different from nash1950bargaining’s (nash1950bargaining) original formalization. In the most standard framework of the bargaining problem, feasible sets are assumed to be convex, and bargaining solutions are formalized as choice functions. By contrast, we do not assume the convexity of feasible sets and allow solutions to select multiple outcomes. The non-convex bargaining problem has been considered in several studies, such as kaneko1980extension, herrero1989nash, conley1991bargaining, zhou1997nash, ok1999revealed, mariotti2000maximal, xu2006alternative; xu2013rationality; xu2019equitable, and peters2012wpo. For surveys, see also xu2020nonconvex and Thomson2022RED. The convexity has been justified through the assumptions that randomization among agreements is allowed and that each player follows the expected utility theory. In the non-convex problem, we can analyze the situations in which these assumptions do not hold. Note that in such cases, scale invariance still has compelling interpretations, even though this axiom has traditionally been justified within the expected utility framework. For a detailed discussion, see xu2006alternative.
A coarse Nash solution can be regarded as the combination of maximizers of multiple choice criteria, as illustrated in the examples we discussed earlier. The idea of conjoining multiple choice criteria is closely related to tadenuma2002efficiency; tadenuma2005egalitarian and houy2009lexicographic, where the choice correspondence is defined through the sequential application of efficiency and fairness criteria in the context of general resource allocation problems. See also manzini2012choice for a general model of the sequential composition of different choice criteria. In contrast to this sequential approach, our choice correspondence takes into account the intersection of multiple criteria. Such intersections typically involve trade-offs, which often result in incompleteness and render the resulting predictions coarse (e.g., sen1973economic; kamaga2018utilitarianism).
@@ A C Proof of Theorem 1
We first show the only-if part. Let $F$ be a solution that satisfies weak Arrow axiom, IIE, efficiency, scale invariance, and continuity. By (2) of Lemma 1, we have known that
@@
Next, we prove the if part. It is easy to prove that the coarse Nash solutions satisfy efficiency. To verify the remaining axioms, let $F^{A}$ be some coarse Nash solution.